one side of a triangle is 2. leass than the base, and one side is 3 in. greater than the base. if the erimeter of the triangle is greater than 25 in., how long is the base?
Answer by mathmandotcom(31)

x=base
x-2=s1
x+3=s2
p>25
p=x+x-2+x+3>25
3x+1>25
3x>24
x>8